Bibliography on Linear Logic

This bibliography is a by-product of the seminar on Linear Logic and Applications held at Carnegie Mellon University in Spring 1995. This seminar was organized by Frank Pfenning, Iliano Cervesato and Carsten Schürmann. Although it concentrates on linear logic and its applications in computer science, a few related references concerning other substructural logics (such as relevance logic or the Lambek calculus) are included.

This, like any bibliography, is incomplete. Any corrections, updates, comments, suggestion, new entries, new URL’s for papers, etc. are very much appreciated. In particular, help us filling the “??”. Please, send feedback to the maintainers.

When compiling this bibliography, we had the chance to access material from Andre Scedrov’s and Harold Schellinx and Anne S. Troelstra’s linear logic bibliographies, as well as from the DIKU types bibliography. We would like to acknowledge their authors.

This document is available in the following formats:

   • BibTeX:  http://www.cs.cmu.edu/~carsten/linearbib/linear.bib
   • Html:  http://www.cs.cmu.edu/~carsten/linearbib/linearbib.html
   • Postscript: http://www.cs.cmu.edu/~carsten/linearbib/linearbib.ps
   • DVI:  http://www.cs.cmu.edu/~carsten/linearbib/linearbib.dvi

This bibliography is currently maintained by Iliano Cervesato (iliano@cs.cmu.edu) with the help of Carsten Schürmann (carsten@cs.cmu.edu) and Frank Pfenning (fp@cs.cmu.edu).

Last modified: May 7, 1998.

[1] Stål Aanderaa and Hermann R. Jervell. Recursive inseparability in linear logic. In E. Börger, G. Jäger, H. K. Bünig, S. Martini, and M. M. Richter, editors, Proceedings of the Sixth Workshop on Computer Science Logic, pages 5–13, San Miniato, Italy, September 1992. Springer-Verlag LNCS 702.

[2] Martin Abadi, George Gonthier, and Jean-Jacques Levy. The geometry of optimal lambda reduction. In Proceedings of the Nineteenth Annual Symposium on Principles of Programming Languages, pages 15–26, Albuquerque, New Mexico, January 1992. ACM Press.

[3] Martin Abadi, George Gonthier, and Jean-Jacques Levy. Linear logic without boxes. In Seventh Annual Symposium on Logic in Computer Science, pages 223–234, Santa Cruz, California, June 1992. IEEE Computer Society Press.

[4] Samson Abramsky. An introduction to ‘On the π-calculus and linear logic’ by Gianluigi Bellin and Philip Scott. Manuscript, 1992.

[5] Samson Abramsky. Computational interpretations of linear logic. Theoretical Computer Science, 111:3–57, 1993.

[6] Samson Abramsky. Interaction categories and typed concurrent programming. In Proceedings of the International Summer School of Marktoberdorf, NATO Advanced Science Institutes, series F, pages ??–?? Springer-Verlag, 1994.

[7] Samson Abramsky and Radha Jagadeesan. New foundations for the geometry of interaction. In Seventh Annual Symposium on Logic in Computer Science, pages 211–222, Santa Cruz, California, June 1992. IEEE Computer Society Press.

[8] Samson Abramsky and Radha Jagadeesan. Games and full completeness for multiplicative linear logic. Journal of Symbolic Logic, 59(2):543–574, 1994.

[9] Samson Abramsky and Radha Jagadeesan. New foundations for the geometry of interaction. Information and Computation, 111(1):53–119, 1994. Extended version of the paper published in the Proceedings of LICS’92.

[10] Samson Abramsky and Steven Vickers. Quantales, observational logic, and process semantics. Mathematical Structures in Computer Science, 3:161–227, 1993.

[11] V. Michele Abrusci. Additional results on intuitionistic linear propositional logic. Technical Report 6, Department of Philosophy, University of Bari, Italy, October 1988.

[12] V. Michele Abrusci. Non-commutative classical linear propositional logic, I. Technical Report 8, Department of Philosophy, University of Bari, Italy, March 1989.

[13] V. Michele Abrusci. A comparison between Lambek syntactic calculus and intuitionistic linear propositional logic. Zeitschrift für Mathematische Logik und Grundlagen der Mathematik, 36:11–15, 1990.

[14] V. Michele Abrusci. Cut elimination theorem for pure non-commutative classical linear propositional logic. Technical Report 12, Department of Philosophy, University of Bari, Italy, September 1990.

[15] V. Michele Abrusci. Non-commutative intuitionistic linear propositional logic. Zeitschrift für Mathematische Logik und Grundlagen der Mathematik, 36:297–318, 1990.

[16] V. Michele Abrusci. Sequent calculus for intuitionistic linear propositional logic. In P. P. Petkov, editor, Mathematical Logic, pages 223–242, New York, London, 1990. Plenum Press. Proceedings of the Summer School and Conference on Mathematical Logic, honourably dedicated to the 90th Anniversary of Arend Heyting (1898–1980), Chaika, Bulgaria, 1988.

[17] V. Michele Abrusci. Lambek syntactic calculus and non-commutative linear logic. In G. Corsi and G. Sambin, editors, Nuovi problemi della logica e della filosofia della scienza, volume II, Viareggio, Italy, January 1991. CLUEB.

[18] V. Michele Abrusci. Phase semantics and sequent calculus for pure non-commutative classical linear propositional logic. Journal of Symbolic Logic, 56(4):1403–1451, December 1991.

[19] V. Michele Abrusci. Non-commutative proof nets. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 271–296. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[20] P. Ageron. Structure des Logiques et Logique des Structures: Logiques, Catégories, Esquisses. PhD thesis, University of Paris VII, January 1991.

[21] David Albrecht, Frank A. Bäuerle, John N. Crossley, and John S. Jeavons. Curry-Howard terms for linear logic. In ??, editor, Logic Colloquium ’94, pages ??–?? ??, 1994.

[22] Vladimir Alexiev. Applications of linear logic to computation: An overview. Technical Report TR93-18, University of Alberta, Edmonton, Alberta T6G 2H1, September 1994.

[23] Gerard Allwein. The Duality of Algebraic and Kripke Models for Linear Logic. PhD thesis, Indiana University, 1992.

[24] Gerard Allwein. A neighborhood model for linear logic’s exponentials. Manuscript, 1992.

[25] Gerard Allwein and Jon M. Dunn. Kripke models for linear logic. Journal of Symbolic Logic, 58(2):514–545, 1993.

[26] S. J. Ambler. First Order Linear Logic in Symmetric Monoidal Closed Categories. PhD thesis, University of Edinburgh, 1991.

[27] Gilles Amiot. Decision problems for second order linear logic without exponentials. Manuscript, 1990.

[28] Jean-Marc Andreoli. Proposal for a Synthesis of Logic and Object-Oriented Programming Paradigms. PhD thesis, University of Paris VI, 1990.

[29] Jean-Marc Andreoli. Logic programming with focusing proofs in linear logic. Journal of Logic and Computation, 2(3):297–347, 1992.

[30] Jean-Marc Andreoli, Tiziana Castagnetti, and Remo Pareschi. Abstract interpretation of linear logic programming. In D. Miller, editor, Proceedings of the International Logic Programming Symposium, pages 295–314, Vancouver, Canada, October 1993. MIT Press.

[31] Jean-Marc Andreoli, Paolo Ciancarini, and Remo Pareschi. Interaction abstract machines. In G. A. Agha, P. Wegner, and A. Yonezawa, editors, Research Directions in Concurrent Object Oriented Programming, pages 257–280. MIT Press, 1993.

[32] Jean-Marc Andreoli, Lone Leth, Remo Pareschi, and Bent Thomsen. True concurrency semantics for a linear logic programming language with broadcast communication. In M.- C. Gaudel and J.-P. Jouannaud, editors, Proceedings of International Joint Conference on Theory and Practice of Software Development, pages 182–198, Orsay, France, April 1993. Springer-Verlag LNCS 668.

[33] Jean-Marc Andreoli and Remo Pareschi. LO and behold! Concurrent structured processes. In Proceedings of OOPSLA’90, pages 44–56, Ottawa, Canada, October 1990. Published as ACM SIGPLAN Notices, vol.25, no.10.

[34] Jean-Marc Andreoli and Remo Pareschi. Communication as fair distribution of knowledge. In Proceedings of OOPSLA’91, pages 212–229, Phoenix, Arizona, October 1991. Published as ACM SIGPLAN Notices, vol.26, no.11.

[35] Jean-Marc Andreoli and Remo Pareschi. Linear objects: Logical processes with built-in inheritance. New Generation Computing, 9:445–473, 1991.

[36] Jean-Marc Andreoli and Remo Pareschi. Logic programming with sequent systems: A linear logic approach. In P. Schröder-Heister, editor, Proceedings of Workshop to Extensions of Logic Programming, Tübingen, 1989, pages 1–30. Springer-Verlag LNAI 475, 1991.

[37] Jean-Marc Andreoli and Remo Pareschi. Associative communication and its optimization via abstract interpretation. Manuscript, August 1992.

[38] D. A. Archangelsky and Mikhail A. Taitslin. Modal linear logic. In A. Nerode and M. Taitslin, editors, Proceedings of the Symposium on Logical Foundations of Computer Science, pages 1–8, Tver, Russia, July 1992. Springer-Verlag LNCS 620.

[39] D. A. Archangelsky and Mikhail A. Taitslin. Linear logic with fixed resources. Journal of Pure and Applied Logic, 67(1-3):3–28, 1994.

[40] J. Arima and H. Sawamura. Reformulation of explanation by linear logic: Toward logic for explanation. In K. P Jantke, editor, 4th International workshop on Algorithmic learning theory, number 744 in LNCS, pages 45–58, Tokyo, Japan, November 1993.

[41] Andrea Asperti. A logic for concurrency. Technical Report ??, Department of Computer Science, University of Pisa, 1987.

[42] Andrea Asperti. Categorical Topics in Computer Science. PhD thesis, University of Pisa, [43] Andrea Asperti. Linear logic, comonads and optimal reductions. Manuscript, to appear in

Fundamenta Informaticae, 1994, special issue devoted to Categories in Computer Science, September 1993.

[44] Andrea Asperti. Causal dependencies in multiplicative linear logic with MIX. Mathematical Structures in Computer Science, to appear.

[45] Andrea Asperti and Giovanna Dore. Yet another correctness criterion for multiplicative linear logic with MIX. Manuscript, November 1993.

[46] Andrea Asperti, Gianluigi Ferrari, and Roberto Gorrieri. Implicative formulae in the ‘proofs as computations’ analogy. In Proceedings of the seventeenth Symposium on Principles of Programming Languages, San Francisco, pages 59–71. ACM Press, January 1990.

[47] Arnon Avron. Relevant entailment: Semantics and formal systems. Journal of Symbolic Logic, 49:334–342, 1984.

[48] Arnon Avron. The Semantics and Proof-Theory of Relevance Logic and Non-Trivial Theories Containing Contradictions. PhD thesis, Tel-Aviv University, 1984.

[49] Arnon Avron. The semantics and proof theory of linear logic. Theoretical Computer Science, 57:161–184, 1988.

[50] Arnon Avron. Simple consequence relations. Information and Computation, 92:105–139, 1991.

[51] Arnon Avron. Some properties of linear logic proved by semantic methods. Journal of Logic and Computation, 4(6):929–938, 1994.

[52] Arnon Avron. Multiplicative conjunction and the algebraic meaning of contraction and weakening. Journal of Symbolic Logic, to appear.

[53] Henry Baker. Lively linear lisp — ‘Look Ma, no garbage!’. ACM SIGPLAN Notices, 27(8):89–98, August 1992.

[54] Henry Baker. NREVERSAL of fortune–the thermodynamics of garbage collection. In Y. Bekkers and J. Cohen, editors, International Workshop on Memory Management, pages 507–524, Saint-Malo, France, September 1992. Springer-Verlag LNCS 637.

[55] Henry Baker. The Boyer benchmark meets linear logic. ACM Lisp Pointers, VI(4):3–10, 1993.

[56] Henry Baker. Sparse polynomials and linear logic. ACM Sigsam Bulletin, 27(4):10–14, December 1993.

[57] Henry Baker. Linear logic and permutation stacks — the forth shall be first. ACM Computer Architecture News, 22(1):34–43, March 1994.

[58] Henry Baker. A “linear logic” quicksort. SIGPLAN Notices, 29(2):13–18, February 1994.

[59] Henry Baker. Minimizing reference count updating with deferred and anchored pointers for functional data structures. ACM SIGPLAN Notices, 29(9):38–43, September 1994.

[60] Henry Baker. ‘Use-once’ variables and linear objects — storage management, reflexion and multi-threading. ACM SIGPLAN Notices, 30(1):45–52, January 1995.

[61] Michael Barr. ⋆-Autonomous Categories. Springer-Verlag LNM 752, 1979.

[62] Michael Barr. Accessible categories and models of linear logic. Journal of Pure and Applied Algebra, 69:219–232, 1990.

[63] Michael Barr. Fuzzy models of linear logic. Preprint, 1991.

[64] Michael Barr. ⋆-Autonomous categories and linear logic. Mathematical Structures in Computer Science, 1:159–178, 1991.

[65] Denis Bechet, Philippe de Groote, and Christian Retoré. A complete axiomatisation for the inclusion of series-parallel partial orders. In Proccedings of the Eighth International Conference on Rewriting Techniques and Applications, Sitges, Spain, 1997.

[66] Giangluigi Bellin. Subnets of proof-nets in multiplicative linear logic with mix. Mathematical Structures in Computer Science, to appear.

[67] Giangluigi Bellin and A. Fleury. Braided proof-nets for multiplicative linear logic with Mix. Manuscript, December 1995.

[68] Gianluigi Bellin. Mechanizing Proof Theory: Resource-Aware Logics and Proof-Transformations to Extract Implicit Information. PhD thesis, Stanford University, 1990.

[69] Gianluigi Bellin. Proof-nets for multiplicative and additive linear logic. Technical Report LFCS-91-161, Deparment of Computer Science, University of Edinburgh, May 1991.

[70] Gianluigi Bellin. Proof nets for classical logic LC. Manuscript, 1992.

[71] Gianluigi Bellin. Proof nets for multiplicative and additive linear logic. Manuscript, May 1993.

[72] Gianluigi Bellin. Empires and kingdoms in DL. Manuscript, March 1994.

[73] Gianluigi Bellin. On proof nets for multiplicative and additive linear logic: Correctness conditions and intuitionistic translations. Manuscript, October 1994.

[74] Gianluigi Bellin. Proof-nets for LC, with an appendix on LG. Manuscript, July 1994.

[75] Gianluigi Bellin. Empires and kingdoms in MLL- and MIX. Mathematical Structures in Computer Science, to appear.

[76] Gianluigi Bellin and Jussi Ketonen. A decision procedure revisited: Notes on direct logic, linear logic and its implementation. Theoretical Computer Science, 95:115–142, 1992.

[77] Gianluigi Bellin and Philip J. Scott. On the π-calculus and linear logic. Manuscript, 1992.

[78] Gianluigi Bellin and J. van de Wiele. Proof nets and typed lambda calculus I: Empires and kingdoms. Manuscript, May 1993.

[79] Gianluigi Bellin and J. van de Wiele. Empires and kingdoms in MLL. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 249–270. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[80] Nuel Belnap Jr. Linear logic displayed. Notre-Dame Journal of Formal Logic, 31:14–25, 1990.

[81] Nick Benton. Strong normalization for the linear term calculus. Journal of Functional Programming, 5(1):65–80, 1995. Also available as Technical Report 305, Computer Laboratory, University of Cambridge, July 1993.

[82] Nick Benton, G. M. Bierman, J. Martin E. Hyland, and Valeria de Paiva. Linear λ-calculus and categorical models revisited. In E. Börger, editor, Proceedings of the Sixth Workshop on Computer Science Logic, pages 61–84. Springer-Verlag LNCS 702, 1992.

[83] Nick Benton, G. M. Bierman, J. Martin E. Hyland, and Valeria de Paiva. Term assignment for intuitionistic linear logic. Technical Report 262, Computer Laboratory, University of Cambridge, August 1992.

[84] Nick Benton, G. M. Bierman, J. Martin E. Hyland, and Valeria de Paiva. A term calculus for intuitionistic linear logic. In M. Bezem and J. F. Groote, editors, Proceedings of the International Conference on Typed Lambda Calculi and Applications, pages 75–90. Springer-Verlag LNCS 664, 1993.

[85] P. Benton. A mixed linear and non-linear logic: Proofs, terms and models (preliminary report). Technical Report 352, University of Cambridge, Computer Laboratory, September 1994.

[86] G. Berry and G. Boudol. The chemical abstract machine. In Conference Record of the Seventeenth Annual Symposium on Principles of Programming Languages, pages 81–94, San Francisco, California, January 1990.

[87] Inge Bethke. Coherence spaces are untopological. Theoretical Computer Science, 85:353– 357, 1991.

[88] Wofgang Bibel, Luis Fariñas del Cerro, B. Fronhöfer, and A. Herzig. Plan generation by linear proofs: On semantics. In German Workshop on Artificial Intelligence - GWAI’89, volume 216 of Informatik-Fachberichte, pages ??–?? Springer-Verlag, 1989.

[89] G. M. Bierman. Type systems, linearity and functional programming languages. Manuscript, December 1991.

[90] G. M. Bierman. On Intuitionistic Linear Logic. PhD thesis, University of Cambridge, 1994.

[91] G. M. Bierman. On intuitionistic linear logic. Technical Report 346, University of Cambridge, Computer Laboratory, August 1994. Revised version of PhD thesis.

[92] G. M. Bierman. What is a categorical model of intuitionistic linear logic? In M. Dezani, editor, Proceedings of Conference on Typed lambda calculus and Applications. Springer-Verlag LNCS 902, April 1995. An earlier version appeared as University of Cambridge Computer Laboratory Technical Report 333, April 1994.

[93] G. M. Bierman. A classical linear lambda calculus. Technical Report 401, University of Cambridge Computer Laboratory, July 1996.

[94] G. M. Bierman. A note on full intuitionistic linear logic. Annals of Pure and Applied Logic, 79(3):281–287, June 1996.

[95] G. M. Bierman. Towards a classical linear lambda calculus. In Proceedings of Tokyo Conference on Linear Logic, volume 3 of Electronic Notes in Computer Science, Tokyo, Japan, 1996. Elsevier. To appear.

[96] G.M. Bierman. Observations on a linear PCF. Technical Report 412, University of Cambridge Computer Laboratory, January 1997.

[97] Andreas Blass. A game semantics for linear logic. Annals of Pure and Applied Logic, 56:183–220, 1992. Special Volume dedicated to the memory of John Myhill.

[98] Andreas Blass. A category arising in linear logic, complexity theory, and set theory. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 61–81. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[99] Richard Blute. Linear Logic, Coherence and Dinaturality. PhD thesis, University of Pennsylvania, 1991.

[100] Richard Blute. Proof nets and coherence theorems. In S. Abramsky, P.-L. Curien, A. M. Pitt, D. H. Pitts, A. Poigné, and D. E. Rydeheard, editors, Category Theory and Computer Science, Springer-Verlag LNCS 530, pages 121–137, Paris, France, September 1991.

[101] Richard Blute. ⋆-autonomous hyperdoctrines and predicate linear logic. Manuscript, 1992.

[102] Richard Blute. Linear logic, coherence and dinaturality. Theoretical Computer Science, 115:3–41, 1993.

[103] Richard Blute. Hopf algebras and linear logic. Mathematical Structures in Computer Science, to appear.

[104] Richard Blute, J. R. B. Cockett, and R. A. G. Seely. Categories for computation in context and unified logic. Manuscript, 1996.

[105] Richard Blute, J. R. B. Cockett, and R. A. G. Seely. ! and ? storage as tensorial strength. Mathematical Structures in Computer Science, to appear.

[106] Richard Blute, J. R. B. Cockett, R. A. G. Seely, and T. H. Trimble. Natural deduction and coherence for weakly distributive categories. Journal of Pure and Applied Algebra, to appear.

[107] Richard Blute, Prakash Panangaden, and R. A. G. Seely. Holomorphic models of exponential types in linear logic. In S. Brookes, M. Main, A. Melton, M. Mislove, and D. Schmidt, editors, Proceedings of the Nineth International Conference on Mathematical Foundations of Programming Semantics, pages 474–512, New Orleans, Louisiana, April 1993.

[108] Richard Blute, Prakash Panangaden, and R. A. G. Seely. Fock space: A model of linear exponential types. Manuscript, revised version of the MFPS paper above, 1994.

[109] Richard Blute and Philip J. Scott. Linear Läuchli semantics. Manuscript, 1995.

[110] Richard Blute and Philip J. Scott. The shuffle Hopf algebra and noncommutative full completeness. Manuscript, 1996.

[111] A. W. Bollen. Relevant logic programming. Journal of Automated Reasoning, 7(4):563– 586, December 1990.

[112] G. Boudol. The lambda-calculus with multiplicities. Manuscript, 1993.

[113] G. Boudol. Some chemical abstract machines. Technical Report BP 93, INRIA-Sophia Antipolis, 1993.

[114] Torben Braüner. The Girard translation extended with recursion. In L. Pacholski and J. Tiuryn, editors, Proceedings of the 1994 Annual Conference of the European Association for Computer Science Logic, pages 31–45, Kazimierz, Poland, September 1994. Springer Verlag, LNCS 933. Short version as Technical Report BRICS-RS-95-13, BRICS, Aarhus, Danemark.

[115] Torben Braüner. A model of intuitionistic affine logic from stable domain theory. In S. Abiteboul and E. Shamir, editors, Proceedings of the 21st International Colloquium on Automata, Languages, and Programming, pages 340–351, Jerusalem, Israel, July 1994. Springer Verlag, LNCS 820. Short version as Technical Report BRICS-RS-94-27, BRICS, Aarhus, Danemark.

[116] Torben Braüner. An Axiomatic Approach to Adequacy. PhD thesis, BRICS, Aarhus, Danemark, November 1996. Also available as Technical Report BRICS-DS-96-4, BRICS, Aarhus, Danemark.

[117] Torben Braüner. A simple adequate categorical model for PCF. In R. Hindley, editor, Proceedings of the 3rd International Converence on Typed Lambda Calculi and Applications, Nancy, France, 1997. To appear.

[118] Torben Braüner. A general adequacy result for a linear functional language. Theoretical Computer Science, To appear.

[119] Torben Braüner and Valeria de Paiva. Cut-elimination for full intuitionistic linear logic. Technical Report BRICS-RS-96-10, BRICS, Aarhus, Danemark, 1996. Also available as Technical Report 395, Computer Laboratory, University of Cambridge.

[120] C. Brown. Relating Petri nets to formulas of linear logic. Technical Report ??, University of Edinburgh, 1989.

[121] C. Brown. Linear Logic and Petri Nets: Categories, Algebra and Proof. PhD thesis, University of Edinburgh, 1990.

[122] C. Brown. Linear logic, Petri nets and quantales. Manuscript, March 1990.

[123] C. Brown and D. Gurr. A categorical linear framework for Petri nets. In Proceedings of the Sixth Annual Symposium on Logic in Computer Science, pages 208–219, Philadelphia, Pennsylvania, June 1990. IEEE Computer Society Press.

[124] C. Brown and D. Gurr. Relations and non-commutative linear logic. Technical Report DAIMI PB-372, Computer Science Department, Aarhus University, November 1991.

[125] C. Brown and D. Gurr. A representation theorem for quantales. Journal of Pure and Applied Algebra, 85:27–42, 1993.

[126] C. Brown, D. Gurr, and Valeria de Paiva. A linear specification language for Petri nets. Technical Report DAIMI PB-363, Computer Science Department, Aarhus University, October 1991.

[127] A. Bucalo. Modalities in linear logic weaker than the exponential ‘of-course’: Algebraic and relational semantics. Manuscript, 1993.

[128] S. Castellani and Paolo Ciancarini. Comparative semantics of LO. Technical Report UBLCS-94-7, University of Bologna, April 1994.

[129] Serenella Cerrito. A linear semantics for allowed logic programs. In Proceedings of the Fifth Symposium on Logic in Computer Science, pages 219–227, Philadelphia, Pennsylvania, June 1990. IEEE Computer Society Press.

[130] Serenella Cerrito. A linear axiomatization of negation as failure. Journal of Logic Programming, 12:1–24, 1992.

[131] Serenella Cerrito. Negation and linear completion. In L. Fariñas del Cerro and M. Penttonen, editors, Intensional Logic for Programming, pages 155–194. Clarendon Press, 1992.

[132] Iliano Cervesato. Lollipops taste of vanilla too. In A. Momigliano and M. Ornaghi, editors, Proof-Theoretical Extensions of Logic Programming, pages 60–66, Santa Margherita Ligure, Italy, June 1994. Post-Conference Workshop for ICLP ’94.

[133] Iliano Cervesato, Joshua S. Hodas, and Frank Pfenning. Efficient resource management for linear logic proof search. In R. Dyckhoff, H. Herre, and P. Schroeder-Heister, editors, Proceedings of the Fifth International Workshop on Extensions of Logic Programming — ELP’96, pages 67–81, Leipzig, Germany, 28–30 March 1996. Springer-Verlag LNAI 1050.

[134] Iliano Cervesato and Frank Pfenning. Linear higher-order pre-unification. In D. Galmiche, editor, Proceedings of the International Workshop on Proof-Search in Type-Theoretic Languages, pages 41–50, New Brunswick, NJ, 30 july 1996.

[135] Iliano Cervesato and Frank Pfenning. A linear logical framework. In E. Clarke, editor, Proceedings of the Eleventh Annual Symposium on Logic in Computer Science — LICS’96, pages 264–275, New Brunswick, New Jersey, 27–30 July 1996. IEEE Computer Society Press. This work also appeared as Preprint 1834 of the Department of Mathematics of Technical University of Darmstadt, Germany.

[136] H. Chau. A proof search system for a modal substructural logic based on labelled deductive systems. In A. Voronkov, editor, Proceedings of the Conference on Logic Programming and Automated Reasoning, pages 64–75. Springer-Verlag LNAI 624, 1992. Extended abstract available as Technical Report DOC 93/1, Imperial College, London.

[137] Jawahar L. Chirimar. Proof Theoretic Approach to Specification Languages. PhD thesis, Department of Computer and Information Science, University of Pennsylvania, 1995.

[138] Jawahar L. Chirimar, Carl A. Gunter, and Jon G. Riecke. Proving memory management invariants for a language based on linear logic. Lisp and Functional Programming, pages 139–150, 1992.

[139] Jawahar L. Chirimar, Carl A. Gunter, and Jon G. Riecke. Reference counting as a computational interpretation of linear logic. Journal of Functional Programming, to appear.

[140] Jawahar L. Chirimar, Carl A. Gunter, and M. van Inwegen. Xpnet: A graphical interface to proof nets with an efficient proof checker. In D. Kapur, editor, Proceedings of the 11th International Conference on Automated Deduction, pages 711–715. Springer-Verlag LNAI 607, June 1992.

[141] Jawahar L. Chirimar and James Lipton. Probability in TBLL: A decision procedure. In E. Börger, G. Jäger, H. K. Bünig, and M. M. Richter, editors, Proceedings of the Fifth Workshop on Computer Science Logic, pages 341–356, Berne, Switzerland, October 1991. Springer-Verlag LNCS 626.

[142] J. R. B. Cockett and R. A. G. Seely. Proof theory for full intuitionistic linear logic, bilinear logic, and MIX categories. Manuscript, 1996.

[143] J. R. B. Cockett and R. A. G. Seely. Weakly distributive categories. Journal of Pure and Applied Algebra, to appear.

[144] M. Coniglio and F. Miraglia. Equality in linear logic. Manuscript, 1995.

[145] M. D’Agostino and Dov M. Gabbay. Fibred tableaux for multi-implication logics. In P. Miglioli, U. Moscato, D. Mundici, and M. Ornaghi, editors, TABLEAUX ’96. Proceedings of the 5th International Workshop on Theorem Proving with Analytic Tableaux and Related Methods, pages 1–35. Springer-Verlag LNAI 1071, 1996.

[146] Marcello D’Agostino and Dov M. Gabbay. A generalization of analytic deduction via labelled deductive systems I: Basic substructural logics. Journal of Automated Reasoning, 13:243–281, 1994.

[147] Marcello D’Agostino, Dov M. Gabbay, and Alessandra Russo. Grafting modalities onto substructural implication logics. To appear in Studia Logica, special issue on “Combining Logics”, n. 4 and 5, 1997, 1997.

[148] Mads Dam. Process-algebraic interpretation of positive linear and relevant logics. Journal of Logic and Computation, 4(6):939–973, 1994.

[149] Vincent Danos. Logique linéaire: Une représentation algébrique du calcul. Gazette des Mathématiciens (Societé Mathématique de France), 41:55–64, 1989.

[150] Vincent Danos. La Logique Linéaire Appliquée à l’étude de Divers Processus de Normalisation (Principalement du λ-Calcul). PhD thesis, University of Paris VII, June 1990.

[151] Vincent Danos. Lambda-calculus and its dynamic algebra I: basic results. Preprint 7, Équipe de Logique Mathématique, University of Paris VII, February 1990.

[152] Vincent Danos, Jean-Baptiste Joinet, and Harold Schellinx. LKT and λμ-calculus. Manuscript, 1993.

[153] Vincent Danos, Jean-Baptiste Joinet, and Harold Schellinx. The structure of exponentials: Uncovering the dynamics of linear logic proofs. In G. Gottlob, A. Leitsch, and D. Mundici, editors, Proceedings of the Third Kurt Gödel Colloquium on Computational Logic and Proof Theory, pages 159–171, Brno, Czech Republic, August 1993. Springer-Verlag LNCS 348.

[154] Vincent Danos, Jean-Baptiste Joinet, and Harold Schellinx. A new deconstructive logic: linear logic. Preprint 52, Équipe de Logique Mathématique, University of Paris VII, 1994.

[155] Vincent Danos, Jean-Baptiste Joinet, and Harold Schellinx. On the linear decoration of intuitionistic derivations. In Archive for Mathematical Logic, volume 33, pages 387–412, 1995.

[156] Vincent Danos, Jean-Baptiste Joinet, and Harold Schellinx. Sequent calculi for second order logic. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 211–224. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[157] Vincent Danos and Laurent Regnier. The structure of multiplicatives. Archive for Mathematical Logic, 28:181–203, 1989.

[158] Vincent Danos and Laurent Regnier. Proof-nets and Hilbert space. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 307–328. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[159] Vincent Danos and Laurent Regnier. Reversible and irreversible computations. Manuscript, 1995.

[160] G. K. Dardzhania. Intuitionistic system without contraction. Polish Academy of Science, Institute of Philosophy and Sociology, Bulletin of the Section of Logic, 6:2–8, 1977.

[161] Philippe de Groote. Linear logic with isabelle : pruning the proof search tree. In A. Bundy, editor, Automated deduction, CADE-12 : 12th International Conference on Automated Deduction, Nancy, France, June/July 1995. Springer-Verlag LNCS 814.

[162] Philippe de Groote. Partially commutative linear logic: Sequent calculus and phase semantics. In V. M. Abrusci and C. Casadio, editors, Proofs and Linguistic Categories, Application of Logic to the Analysis and Implementation of Natural Language, pages 199– 208, Roma, 1996. Cooperativa Libraria Universitaria Editrice Bologna.

[163] Valeria de Paiva. The Dialectica categories. In J. W. Gray and A. Scedrov, editors, Proceedings of the Conference on Categories in Computer Science and Logic, pages 47–62. American Mathematical Society, 1989. AMS-IMS-SIAM Joint Summer Research Conference, June 14–20, 1987, Boulder, Colorado; Contemporary Mathematics Volume 92.

[164] Valeria de Paiva. A Dialectica-like model of linear logic. In P. Dybjer, A. M. Pitts, D. H. Pitt, A. Poigné, and D. E. Rydeheard, editors, Proceeding of the Conference Category Theory and Computer Science, pages 341–356, Manchester, United Kingdom, September 1989. Springer-Verlag LNCS 389.

[165] Valeria de Paiva. Categorical multirelations, linear logic and Petri nets. Manuscript, May 1991.

[166] Valeria de Paiva. The Dialectica categories. Technical Report 213, University of Cambridge, Computer Laboratory, January 1991. Slightly revised version of PhD thesis.

[167] Akim Demaille. Yet another mixed intuitionistic linear logic. Technical Report 98 INF 002, École Nationale Supérieure des Télécommunications, 1997. Available on

http://www.inf.enst.fr/~demaille/papers.

[168] Akim Demaille. Yet another mixed linear logic. Technical Report 98 INF 001, École Nationale Supérieure des Télécommunications, 1997. Available on

http://www.inf.enst.fr/~demaille/papers.

[169] Akim Demaille. Still other mixed linear logics, first version. Technical Report 98 D 002, École Nationale Supérieure des Télécommunications, 1998.

[170] Akim Demaille. Still other mixed linear logics, second version. Technical report, École Nationale Supérieure des Télécommunications, 1998. Available on

http://www.inf.enst.fr/~demaille/papers.

[171] Kostas Došen. The first axiomatization of relevant logic. Journal of Philosophical Logic, 21:337–354, 1992.

[172] Kostas Došen. Modal translations in substructural logics. Journal of Philosophical Logic, 21:283–336, 1992.

[173] Kostas Došen. Non-modal classical linear predicate logic is a fragment of intuitionistic linear predicate logic. Theoretical Computer Science, 102:207–214, 1992. Also in Colloquium on Modal Logic, a selection of papers presented at the Seminar on Intensional Logic, October 1990 - May 1991, M. de Rijke editor.

[174] Jon M. Dunn. Relevance logic and entailment. In D. Gabbay and F. Günter, editors, Handbook of Philosophical Logic, volume III, pages 117–224. Reidel Publication Company, 1986.

[175] Roy Dyckhoff. Contraction-free sequent calculi for intuitionistic logic. Journal of Symbolic Logic, 57(3):795–807, September 1992.

[176] T. Ehrhard. Hypercoherence: A strongly stable model of linear logic. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 83–108. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[177] M. Elvang-Gøransson, Hermann R. Jervell, E. Monteiro, and A. Waaler. Propositional linear logic: Wiring semantics. Information Processing Letters, May 1991. To appear??

[178] Martin Emms. An undecidibility result for polymorphic lambek calculus. Manuscript, 1995.

[179] U. Engberg and G. Winskel. Petri nets as models of linear logic. In A. Arnold, editor, Proceedings of Colloquium on Trees in Algebra and Programming, pages 147–161, Copenhagen, Denmark, 1990. Springer-Verlag LNCS 389.

[180] U. Engberg and G. Winskel. Completeness results for linear logic on Petri nets. In A. Borzyszkowski and S. Sokolowski, editors, Proceedings of the Conference on Mathematical Foundations of Computer Science, pages 442–452, Gdańsk, Poland, 1993. Springer-Verlag LNCS 711. Full version is DAIMI PB, January 1993.

[181] Gianluigi Ferrari and U. Montanari. Typed additive concurrency. Unpublished, 1994.

[182] Andrzej Filinski. Linear continuations. In Conference Record of the Nineteenth Annual Symposium on Principles of Programming Languages, pages 27–38, Albuquerque, New Mexico, January 1992. ACM Press.

[183] A. Fleury and Christian Rétoré. Logique linéaire: Réseaux du fragment multiplicatif avec éléments neutres. Preprint, University of Paris VII, April 1988.

[184] A. Fleury and Christian Rétoré. The MIX rule: A proof-net syntax for multiplicatives, their units and (a x b) -o (a @ b). Preprint 11, Équipe de Logique Mathématique, University of Paris VII, 1990.

[185] C. Fouqueré and J. Vauzeilles. Linear logic and exceptions. Journal of Logic and Computation, 4(6):860–875, 1994.

[186] C. Fouqueré and J. Vauzeilles. Inheritance with exceptions. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 167–196. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[187] B. Fronhöfer. Linearity and plan generation. New Generation Computing, 5:213–225, 1987.

[188] B. Fronhöfer. Linear proofs and linear logic. In D. Pearce and G. Wagner, editors, Logic in AI: European Workshop JELIA’92, pages 106–125, Berlin, Germany, September 1992. Springer-Verlag LNAI 633.

[189] Dov M. Gabbay. Algorithmic proofs with diminishing resources, part I. In E. Börger, editor, Proceedings of the Fourth Workshop on Computer Science Logic, pages 156–173, Heidelberg, Germany, October 1990. Springer-Verlag LNCS 533.

[190] Dov M. Gabbay and Ruy J. G. B. de Queiroz. Extending the Curry-Howard interpretation to linear, relevant and other resource logics. Journal of Symbolic Logic, 57(4):1319–1365, December 1992.

[191] Jean Gallier. Constructive logics. Part II: Linear logic and proof nets. Research Report PR2-RR-9, Digital Equipment Corporation, Paris, 1991.

[192] Didier Galmiche. Canonical proofs for linear logic programming frameworks. In A. Momigliano and M. Ornaghi, editors, Proof-Theoretical Extensions of Logic Programming, pages 2–10, Santa Margherita Ligure, Italy, June 1994. Post-Conference Workshop for ICLP’94.

[193] Didier Galmiche and E. Boudinet. Proof search for programming in intuitionistic linear logic. In D. Galmiche and L. Wallen, editors, CADE-12 Workshop on Proof Search in Type-Theoretic Languages, pages 24–30, Nancy, France, June 1994.

[194] Didier Galmiche and Guy Perrier. Automated deduction in additive and multiplicative linear logic. In A. Nerode and M. Taitslin, editors, Proceedings of the Symposium on Logical Foundations of Computer Science, pages 151–162, Tver, Russia, July 1992. Springer-Verlag LNCS 620.

[195] Didier Galmiche and Guy Perrier. A procedure for automatic proof nets construction. In A. Voronkov, editor, Proceedings of the International Conference on Logic Programming and Automated Reasoning, pages 42–53, St. Petersburg, Russia, 1992. Springer-Verlag LNAI 624.

[196] Didier Galmiche and Guy Perrier. Foundations of proof search strategies design in linear logic. In Symposium on Logical Foundations of Computer Science, pages 101–113, St. Petersburg, Russia, 1994. Springer-Verlag LNCS 813. Also available as Technical Report CRIN 94-R-112 from the Centre di Recherche en Informatique de Nancy.

[197] Didier Galmiche and Guy Perrier. On proof normalisation in linear logic. Theoretical Computer Science, 135(1):67–110, 1994. Also available as Technical Report CRIN 94-R-113 from the Centre di Recherche en Informatique de Nancy.

[198] Simon Gay. Linear Types for Communicating Processes. PhD thesis, University of London, 1994.

[199] Simon Gay and Rajagopal Nagarajan. A typed calculus of synchronous processes. In D. Kozen, editor, Proceedings of the Tenth Annual Symposium on Logic in Computer Science, San Diego, California, June 1995. To appear.

[200] V. Gehlot. A proof-theoretic approach to semantics of concurrency. PhD thesis, University of Pennsylvania, 1991.

[201] V. Gehlot and Carl A. Gunter. Nets as tensor theories. In G. De Michelis, editor, Proceedings of the tenth International Conference on Application and Theory of Petri Nets, pages 174–191, Bonn, Germany, 1989. Extended and revised version available as Technical Report MS-CIS-89-68, University of Pennsylvania.

[202] V. Gehlot and Carl A. Gunter. Normal process representatives. In Proceedings of Fifth Symposium on Logic in Computer Science, pages 200–207, Philadelphia, Pennsylvania, June 1990.

[203] S. Giambrone. Gentzen System and Decision Procedures for Relevant Logic. PhD thesis, Australian National University, Canberra, 1983.

[204] Jean-Yves Girard. Linear logic. Theoretical Computer Science, 50:1–102, 1987.

[205] Jean-Yves Girard. Linear logic and parallelism. In M. Venturini Zilli, editor, Mathematical Models for the Semantics of Parallelism, pages 166–182. Springer-Verlag LNCS 280, 1987. Advanced School, Rome, September 1986.

[206] Jean-Yves Girard. Multiplicatives. In G. Lolli, editor, Logic and Computer Science: New Trends and Applications, pages 11–34, 1987. Rendiconti del Seminario Matematico dell’Università e Politecnico di Torino.

[207] Jean-Yves Girard. Quantifiers in linear logic. In Temi e prospettivi della logica e della filosofia della scienza contemporanee, volume I, pages ??–??, Cesena, Italy, 1987. CLUEB.

[208] Jean-Yves Girard. Geometry of interaction I: Interpretation of system F. In C. Bonotto, R. Ferro, S. Valentini, and A. Zanardo, editors, Logic Colloquium ’88, pages 221–260. North-Holland, 1989.

[209] Jean-Yves Girard. Towards a geometry of interaction. In J. W. Gray and A. Scedrov, editors, Categories in Computer Science and Logic, pages 69–108. American Mathematical Society, 1989. Proceedings of the AMS-IMS-SIAM Joint Summer Research Conference, June 14–20, 1987, Boulder, Colorado; Contemporary Mathematics Volume 92.

[210] Jean-Yves Girard. Geometry of interaction II: Deadlock-free algorithms. In P. Martin-Löf and G. Mints, editors, COLOG-88, pages 76–93. Springer-Verlag LNCS 417, 1990.

[211] Jean-Yves Girard. La logique linéaire. Pour La Science, Edition Française de ‘Scientific American’, 150:74–85, April 1990.

[212] Jean-Yves Girard. A new constructive logic: Classical logic. Mathematical Structures in Computer Science, 1:255–296, 1991.

[213] Jean-Yves Girard. Quantifiers in linear logic II. In G. Corsi and G. Sambin, editors, Nuovi problemi della logica e della filosofia della scienza, volume II, pages ??–??, Bologna, Italy, 1991. CLUEB. Proceedings of the conference with the same name, Viareggio, Italy, January 1990.

[214] Jean-Yves Girard. Logic and exceptions: A few remarks. Journal of Logic and Computation, 2(2):111–118, 1992. Also available as Prépublication 18, Équipe de Logique Mathématique, University of Paris VII, December 1990.

[215] Jean-Yves Girard. Linear logic: A survey. In L. F. Bauer, W. Brauer, and H. Schwichtenberg, editors, Proceedings of the International Summer School of Marktoberdorf, NATO Advanced Science Institutes, series F94, pages 63–112. Springer-Verlag, 1993. Also in P. De Groote editor, The Curry-Howard Isomorphism, pages 193–255, Département de Philosophie, Université Catholique de Louvain, Cahiers du Centre de Logique 8, Academia Press.

[216] Jean-Yves Girard. On the unity of logic. Annals of Pure and Applied Logic, 59:201–217, 1993.

[217] Jean-Yves Girard. Geometry of interaction III: The general case. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 329–389. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[218] Jean-Yves Girard. Linear logic: Its syntax and semantics. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 1–42. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[219] Jean-Yves Girard. On denotational completeness. Manuscript, 1995.

[220] Jean-Yves Girard. Coherent banach spaces : a continuous denotational semantics. unpublished manuscript, 1996.

[221] Jean-Yves Girard. Constructivit : vers une dualit moniste. unpublished manuscript, 1996.

[222] Jean-Yves Girard. Proof-nets: The parallel syntax for proof-theory. In P. Agliano and A. Ursini, editors, Logic and Algebra, pages ??–?? Marcel Dekker, New York, 1996.

[223] Jean-Yves Girard. Titres et travaux. unpublished manuscript, 1996.

[224] Jean-Yves Girard. Du pourquoi au comment : la thorie de la dmonstration de 1950 nos jours. unpublished manuscript, 1997.

[225] Jean-Yves Girard. Light linear logic. Information and Computation, 143, 1998.

[226] Jean-Yves Girard. On the meaning of logical rules I: syntax vs. semantics. unpublished manuscript, 1998.

[227] Jean-Yves Girard. On the meaning of logical rules II: multiplicatives and additives. unpublished manuscript, 1998.

[228] Jean-Yves Girard and Yves Lafont. Linear logic and lazy computation. In H. Ehrig, R. Kowalski, G. Levi, and U. Montanari, editors, Proceedings of the International Joint Conference on Theory and Practice of Software Development, volume 2, pages 52–66, Pisa, Italy, March 1987. Springer-Verlag LNCS 250.

[229] Jean-Yves Girard, Yves Lafont, and Laurent Regnier, editors. Geometry of Interaction III : accommodating the additives, number 222 in Advances in Linear Logic, London Mathematical Society Lecture Notes Series. Cambridge University Press, 1995.

[230] Jean-Yves Girard, Yves Lafont, and Laurent Regnier, editors. Linear Logic, its syntax and semantics, number 222 in Advances in Linear Logic, London Mathematical Society Lecture Notes Series. Cambridge University Press, 1995.

[231] Jean-Yves Girard, Yves Lafont, and P. Taylor. Proofs and Types. Cambridge Tracts in Theoretical Computer Science 7. Cambridge University Press, 1988.

[232] Jean-Yves Girard, Andre Scedrov, and Philip J. Scott. Bounded linear logic. Logic & Computation 36, MS-CIS-91-59, Department of Computer and Information Science, School of Engineering and Applied Science, University of Pennsylvania, August 1991.

[233] Jean-Yves Girard, Andre Scedrov, and Philip J. Scott. Bounded linear logic: A modular approach to polynomial time computability. Theoretical Computer Science, 97:1–66, 1992. Extended abstract in Feasible Mathematics, S. R. Buss and P. J. Scott editors, Proceedings of the MCI Workshop, Ithaca, NY, June 1989, Birkhauser, Boston, pp. 195–209.

[234] Jean-Yves Girard, Andre Scedrov, and Philip J. Scott. Normal forms and cut-free proofs as natural transformations. In Y. Moschovakis, editor, Logic from Computer Science, volume 21, pages 217–241. Springer-Verlag, MSRI publications edition, 1992.

[235] R. Van Glabbeek and Gordon Plotkin. Configuration structures. In Tenth Annual Symposium on Logic in Computer Science, pages 199–209, San Diego, California, June 1995. IEEE Computer Society.

[236] V. N. Grishin. Predicate and set-theoretic calculi based on logic without contractions. Math. USSR. Izvestija, 18:41–59, 1981.

[237] G. Große, S. Hölldobler, and J. Schneeberger. Linear deductive planning. Technical Report AIDA-92-08, Intellektik, Informatik, Techniche Hochschule Darmstadt, 1992.

[238] Vineet Gupta. Concurrent Kripke structures. In Proceedings of the North American Process Algebra Workshop, Cornell CS-TR-93-1369, August 1993.

[239] Vineet Gupta. Chu Spaces: A Model of Concurrency. PhD thesis, Stanford University, September 1994.

[240] Vineet Gupta and Vaughan R. Pratt. Gates accept concurrent behavior. In Proceedings of the 34th Annual IEEE Symposium on Foundations of Computer Science, pages 62–71, November 1993.

[241] Juan C. Guzmán and Paul Hudak. Single-threaded polymorphic lambda calculus. In Proceedings of the Fifth Symposium on Logic in Computer Science, pages 333–343, Philadelphia, Pennsylvania, June 1990. IEEE Computer Society Press.

[242] James Harland. A proof-theoretic analysis of goal-directed provability. Journal of Logic and Computation, 4(1):69–88, 1994.

[243] James Harland and David Pym. The uniform proof-theoretic foundation of linear logic programming. Technical Report ECS-LFCS-90-124, Laboratory for the Foundations of Computer Science, University of Edinburgh, November 1990.

[244] James Harland and David Pym. The uniform proof-theoretic foundation of linear logic programming. In V. Saraswat and K. Ueda, editors, Proceedings of the International Logic Programming Symposium, pages 304–318, San Diego, California, October 1991.

[245] James Harland and David Pym. On resolution in fragments of classical linear logic. In Proceedings of the Russian Conference on Logic Programming and Automated Reasoning, pages 30–41. Springer-Verlag LNAI 624, July 1992.

[246] James Harland and David Pym. A note on the implementation and applications of linear logic programming languages. In G. Gupta, editor, Proceedings of the Seventeenth Annual Computer Science Conference, Christchurch, pages 647–658, January 1994.

[247] James Harland and David Pym. A uniform proof-theoretic investigation of linear logic programming. Journal of Logic and Computation, 4(2):175–207, April 1994.

[248] James Harland, David Pym, and Michael Winikoff. Programming in Lygon: An overview. In M. Wirsing and M. Nivat, editors, Algebraic Methodology and Software Technology, pages 391–405, Munich, Germany, July 1996. Springer-Verlag LNCS 1101.

[249] James Harland and Michael Winikoff. Deterministic resource management for the linear logic programming language Lygon. Technical Report TR 94/23, Melbourne University, Department of Computer Science, 1994.

[250] James Harland and Michael Winikoff. Implementation and development issues for the linear logic programming language Lygon. In Proceedings of the Eighteenth Australian Computer Science Conference, pages 563–572, Adelaide, Australia, February 1995. Also available as Technical Report TR 95/6, Melbourne University, Department of Computer Science.

[251] James Harland and Michael Winikoff. Implementing the linear logic programming language Lygon. In J. Lloyd, editor, Proceedings of the 1995 International Logic Programming Symposium, pages 66–80, Portland, Oregon, 1995.

[252] James Harland and Michael Winikoff. Some applications of the linear logic programming language Lygon. In Proceedings of the Nineteenth Australasian Computer Science Conference, pages 262–271, Melbourne, Australia, 1996.

[253] Wim H. Hesselink. Axioms and models of linear logic. Formal Aspects of Computing, 2:139–166, 1990.

[254] J. R. Hindley. BCK-combinators and linear λ-terms have types. Theoretical Computer Science, 64:97–105, 1989.

[255] J. R. Hindley. BCK- and BCI-logics, condensed detachment and the 2-property: A summary. Manuscript, September 1990.

[256] Joshua S. Hodas. Lolli: An extension of λprolog with linear context management. In D. Miller, editor, Workshop on the λProlog Programming Language, pages 159–168, Philadelphia, Pennsylvania, August 1992.

[257] Joshua S. Hodas. Specifying filler-gap dependency parsers in a linear-logic programming language. In K. Apt, editor, Proceedings of the Joint International Conference and Symposium on Logic Programming, pages 622–636, Washington, DC, November 1992.

[258] Joshua S. Hodas. Logic programming with multiple context management schemes. In R. Dyckhoff, editor, Fourth International Workshop on Extensions of Logic Programming, pages 171–182, St. Andrews, United Kingdom, March 1993. Springer-Verlag LNCS 360.

[259] Joshua S. Hodas. Logic Programming in Intuitionistic Linear Logic: Theory, Design and Implementation. PhD thesis, University of Pennsylvania, Department of Computer and Information Science, 1994.

[260] Joshua S. Hodas and Dale Miller. Logic programming in a fragment of intuitionistic linear logic. Information and Computation, 110(2):327–365, 1994. Extended abstract in the Proceedings of the Sixth Annual Symposium on Logic in Computer Science, Amsterdam, July 15–18, 1991.

[261] Joshua S. Hodas and Jeffrey Polakow. Forum as a logic programming language: Preliminary results and observations. In M. Okada, editor, Proceedings of the Linear Logic ’96 Meeting, volume 3, Tokyo, Japan, 1996. Elsevier Electronic Notes in Theoretical Computer Science.

[262] S. Holmström. Linear functional programming. In Proceedings of the Workshop on the Implementation of Lazy Functional Languages, pages 13–32, Aspanäes, Sweden, September 1988. Also available as Technical Report 53, Programming Methodoloy Group, Chalmers University of Technology, S-412 96, Göteborg, Sweden.

[263] S. Holmström. Quicksort in a linear functional language. Technical Report PMG Memo 65, Chalmers University of Technology, Göteborg, Sweden, January 1989.

[264] K. Homann and Carsten Schürmann. A combined proof search strategy for linear logic. Manuscript, 1995.

[265] R. Hoofman. Non-Stable Models of Linear Logic. PhD thesis, Utrecht University, April 1992.

[266] R. Hoofman. Non-stable models of linear logic. In A. Nerode and M. Taitslin, editors, Proceedings of the Symposium on Logical Foundations of Computer Science, pages 210– 220, Tver, Russia, July 1992. Springer-Verlag LNCS 620.

[267] R. Hori, Hiroakira Ono, and Harold Schellinx. Extending intuitionistic linear logic with knotted structural rules. Notre-Dame Journal of Formal Logic, 35(2):219–242, 1994.

[268] B. Hösli. Die Abschwächungsregel in der Aussagenlogik. Manuscript, December 1990.

[269] T. Hosoi and Hiroakira Ono. Intermediate propositional logics (a survey). Journal of Tsuda College, 5:67–82, 1973.

[270] J. Martin E. Hyland and Valeria de Paiva. Lineales. Manuscript, February 1991.

[271] J. Martin E. Hyland and Valeria de Paiva. Full intuitionistic linear logic (extended abstract). Annals of Pure and Applied Logic, 64(3):273–291, 1993.

[272] Bart Jacobs. Conventional and linear types in a logic of coalgebras. Manuscript, April 1993.

[273] Bart Jacobs. Semantics of weakening and contraction. Annals of Pure and Applied Logic, 69:73–106, 1994.

[274] Éric Jacopin. Classical AI planning as theorem proving: The case of a fragment of linear logic. In AAAI Fall Symposium on Automated Deduction in Nonstandard Logics, pages 62–66, Palo Alto, California, 1993. AAAI Press Publications.

[275] J-.B. Joinet. Une preuve combinatoire de forte normalisation (du fragment multiplicatif et exponentiel) des réseaux de preuve pour la logique linéaire (pn1). Preprint 35, Équipe de Logique Mathématique, University of Paris VII, 1992.

[276] J-.B. Joinet. Étude de la Normalisation du Calcul des Séquents Classique à Travers la Logique Linéaire. PhD thesis, University of Paris VII, 1993.

[277] P. Jumpertz. Linear logic and real closed fields: A way to handle situations dynamically. Manuscript, 1994.

[278] Max I. Kanovich. The Horn fragment of linear logic is NP-complete. ITLI Prepublication Series X-91-14, University of Amsterdam, 1991.

[279] Max I. Kanovich. The multiplicative fragment of linear logic is NP-complete. ITLI Prepublication Series X-91-13, University of Amsterdam, 1991.

[280] Max I. Kanovich. Horn programming in linear logic is NP-complete. In Seventh Annual Symposium on Logic in Computer Science, pages 200–210, Santa Cruz, California, June 1992. IEEE Computer Society Press.

[281] Max I. Kanovich. The complexity of the Horn fragment of linear logic. Annals of Pure and Applied Logic, 69:195–241, 1994.

[282] Max I. Kanovich. The independent basis of neutral formulas in linear logic. Preprint 94-08, Laboratoire de Mathématiques Discrètes, University of Marseille, 1994.

[283] Max I. Kanovich. Linear logic as a logic of computations. Annals of Pure and Applied Logic, 67(1-3):183–212, 1994. Also in Logic at Tver ’92, Sokal, Russia, July 1992.

[284] Max I. Kanovich. Petri nets, Horn programs, linear logic, and vector games. In M. Hagiya and J. Mitchell, editors, Proceedings of the International Symposium Theoretical Aspects of Computer Software TACS’94, pages 642–666, Sendai, Japan, April 1994. Springer-Verlag LNCS 789.

[285] Max I. Kanovich. Simulating linear logic with 1-linear logic. Preprint 94-02, Laboratoire de Mathématiques Discrètes, University of Marseille, 1994.

[286] Max I. Kanovich. The complexity of neutrals in linear logic. In D. Kozen, editor, Tenth Annual IEEE Symposium on Logic in Computer Science, pages 486–495, San Diego, California, June 1995.

[287] Max I. Kanovich. The direct simulation of Minsky machines in linear logic. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 123–145. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[288] Max I. Kanovich, M. Okada, and A. Scedrov. Phase semantics for light linear logic. unpublished manuscript.

[289] Naoki Kobayashi. Concurrent Linear Logic Programming. PhD thesis, Department of Information Science, University of Tokyo, April 1996.

[290] Naoki Kobayashi, Benjamin C. Pierce, and David N. Turner. Linearity and the pi-calculus. In Proceedings of the 23rd ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages — POPL’96, number 21–24, pages 358–371, St. Petersburg Beach, Florida, January 1996.

[291] Naoki Kobayashi and Akinori Yonezawa. ACL — A concurrent linear logic programming paradigm. In D. Miller, editor, Proceedings of the 1993 International Logic Programming Symposium, pages 279–294, Vancouver, Canada, October 1993. MIT Press.

[292] Naoki Kobayashi and Akinori Yonezawa. Reasoning on actions and change in linear logic programming. Technical Report ??, University of Tokyo, 1993.

[293] Naoki Kobayashi and Akinori Yonezawa. Asynchronous communication model based on linear logic. Formal Aspects of Computing, 3:279–294, 1994. Short version appeared in Joint International Conference and Symposium on Logic Programming, Washington, DC, November 1992, Workshop on Linear Logic and Logic Programming.

[294] Naoki Kobayashi and Akinori Yonezawa. Typed higher-order concurrent linear logic programming. Technical Report 94-12, University of Tokyo, 1994.

[295] Yuichi Komori and Hiroakira Ono. Logics without the contraction rule. Journal of Symbolic Logic, 50(1):169–201, March 1985.

[296] Alexei P. Kopylov. Decidability of linear affine logic. In D. Kozen, editor, Tenth Annual IEEE Symposium on Logic in Computer Science, pages 496–504, San Diego, California, June 1995.

[297] Christoph Kreitz, Heiko Mantel, Jens Otten, and Stephan Schmitt. Connection-based proof construction in linear logic. In W. McCune, editor, Proceedings of the 14th International Conference on Automated Deduction, pages 207–221. Springer-Verlag LNAI 1249, 1997.

[298] Yves Lafont. Introduction to linear logic. Lecture notes for the Summer School on Constructive Logics and Category Theory, 1988.

[299] Yves Lafont. The linear abstract machine. Theoretical Computer Science, 59:157–180, 1988. Some corrections in volume 62 (1988), pp. 327–328.

[300] Yves Lafont. Logiques, Categories et Machines. Implantation de Langages de Programmation guidé par la Logique Catégorique. PhD thesis, University of Paris VII, January 1988.

[301] Yves Lafont. Functional programming and linear logic. Lecture notes for the Summer School on Functional Programming and Constructive Logic, Glasgow, United Kingdom, [302] Yves Lafont. Interaction nets. In Seventeenth Annual Symposium on Principles of Pro-

gramming Languages, pages 95–108, San Francisco, California, 1990. ACM Press.

[303] Yves Lafont. The finite model property for various fragments of linear logic. Manuscript, 1995.

[304] Yves Lafont. From proof nets to interaction nets. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 225–247. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[305] Yves Lafont. Interaction combinators. Manuscript, 1995.

[306] Yves Lafont. The undecidability of second order linear logic without exponentials. Journal of Symbolic Logic, to appear. Also available as preprint 95-06, Laboratoire de Mathématiques discrètes, University of Marseille.

[307] Yves Lafont and Andre Scedrov. The undecidability of second order multiplicative linear logic. Preprint 95-17, Laboratoire de Mathématiques discrètes, University of Marseille, 1995.

[308] Yves Lafont and T. Streicher. Games semantics for linear logic. In Sixth Symposium on Logic in Computer Science, pages 43–50. IEEE Computer Society Press, July 1991. Amsterdam, The Netherlands.

[309] François Lamarche. Games semantics for full propositional linear logic. In D. Kozen, editor, Tenth Annual IEEE Symposium on Logic in Computer Science, pages 464–473, San Diego, California, June 1995.

[310] Joachim Lambek. From categorical grammars to bilinear logic. Manuscript, 1991.

[311] Joachim Lambek. Bilinear logic in algebra and linguistics. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 43–59. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[312] Jean Leneutre. Linear logic: Investigation into the distributivity properties of multiplicatives wrt. additives. Technical report, cole Nationale Suprieure des Tlcommunications, 1998.

[313] J. Lilius. High-level nets and linear logic. In K. Jensen, editor, Proceedings of the International Conference on Applications and Theory of Petri Nets, pages 310–327, Sheffield, United Kingdom, June 1992. Springer-Verlag LNCS 616.

[314] Patrick Lincoln. Computational Aspects of Linear Logic. PhD thesis, Stanford Univeristy, 1992.

[315] Patrick Lincoln. Linear logic. ACM SIGACT Notices, 23(2):29–37, Spring 1992.

[316] Patrick Lincoln. Deciding provability of linear logic formulas. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 109–122. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[317] Patrick Lincoln and John Mitchell. Operational aspects of linear lambda calculus. In Seventh Annual Symposium on Logic in Computer Science, pages 235–246, Santa Cruz, California, June 1992. IEEE Computer Society Press.

[318] Patrick Lincoln, John Mitchell, and Andre Scedrov. Stochastic interaction and linear logic. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 147– 166. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[319] Patrick Lincoln, John Mitchell, and Andre Scedrov. Linear logic proof games and optimization. Manuscript, January 1996.

[320] Patrick Lincoln, John Mitchell, Andre Scedrov, and Natarajan Shankar. Decision problems for propositional linear logic. Annals of Pure and Applied Logic, 56:239–311, April 1992. Also in the Proceedings of the 31th Annual Symposium on Foundations of Computer Science, St Louis, Missouri, October 1990, IEEE Computer Society Press. Also available as Technical Report SRI-CSL-90-08 from SRI International, Computer Science Laboratory.

[321] Patrick Lincoln and Vijay Saraswat. Higher-order, linear, concurrent constraint programming. Manuscript, January 1993.

[322] Patrick Lincoln and Andre Scedrov. First order linear logic without modalities is NEXPTIME-hard. Theoretical Computer Science, 135(1):139–154, 1994.

[323] Patrick Lincoln, Andre Scedrov, and Natarajan Shankar. Linearizing intuitionistic implication. Annals of Pure and Applied Logic, 60:151–177, 1993. Preliminary version in Sixth Annual Symposium on Logic in Computer Science, pages 51–62, Amsterdam, The Netherlands, July 1991.

[324] Patrick Lincoln, Andre Scedrov, and Natarajan Shankar. Decision problems for second order linear logic. In D. Kozen, editor, Tenth Annual IEEE Symposium on Logic in Computer Science, pages 476–485, San Diego, California, June 1995.

[325] Patrick Lincoln and Natarajan Shankar. The complexity of cut elimination in a fragment of linear logic. Manuscript, 1990.

[326] Patrick Lincoln and Natarajan Shankar. Proof search in first-order linear logic and other cut-free sequent calculi. In S. Abramsky, editor, Ninth Annual Symposium on Logic in Computer Science, pages 282–291, Paris, France, 1994. IEEE Computer Society Press.

[327] Patrick Lincoln and Timothy Winkler. Constant-only multiplicative linear logic is NPcomplete. Theoretical Computer Science, 135:155–169, 1994.

[328] Ralph Loader. Linear logic, totality and full completeness. In S. Abramsky, editor, Ninth Annual Symposium on Logic in Computer Science, pages 292–298, Paris, France, 1994. IEEE Computer Society Press.

[329] Ian Mackie. Lilac — a functional programming language based on linear logic. Master’s thesis, Imperial College, London, 1991.

[330] Ian Mackie. λ-calculus, linear logic and the π-calculus: A survey. Manuscript, February 1993.

[331] Ian Mackie. Lilac — a functional programming language based on linear logic. Journal of Functional Programming, 4(4):395–433, 1993.

[332] Pasquale Malacaria and Laurent Regnier. The weak nilpotency of λ-calculus. Manuscript, May 1990.

[333] Pasquale Malacaria and Laurent Regnier. Some results on the interpretation of λ-calculus in operator algebras. In Sixth Symposium on Logic in Computer Science, pages 63–72. IEEE Computer Society Press, July 1991. Amsterdam, The Netherlands.

[334] N. Martí-Oliet and J. Meseguer. From Petri nets to linear logic. In P. Dybjer, A. M. Pitts, D. H. Pitt, A. Poigné, and D. E. Rydeheard, editors, Proceedings of the Conference on Category Theory and Computer Science, Springer-Verlag LNCS 389, pages 313–340, Manchester, United Kingdom, September 1989.

[335] N. Martí-Oliet and J. Meseguer. Duality in closed and linear categories. Technical Report SRI-CSL-90-01, SRI International, Computer Science Laboratory, February 1990.

[336] N. Martí-Oliet and J. Meseguer. An algebraic axiomatization of linear logic models. In G. M. Reed, A. W. Roscoe, and R. F. Wachter, editors, Topology and Category Theory in Computer Science, pages 335–357. Clarendon Press, Oxford, 1991. Proceedings of the Oxford Topology Symposium, June 1989.

[337] N. Martí-Oliet and J. Meseguer. From Petri nets to linear logic. Mathematical Structures in Computer Science, 1:66–101, 1991. Revised version of paper in LNCS 389.

[338] N. Martí-Oliet and J. Meseguer. From Petri nets to linear logic through categories: A survey. Journal on Foundations of Computer Science, 2(4):297–399, December 1991.

[339] Simone Martini and Andrea Masini. An exponential-free interpretation of classical logic into linear logic. Manuscript, 1992.

[340] Simone Martini and Andrea Masini. A modal view of linear logic. Journal of Symbolic Logic, 59(3):888–899, 1994.

[341] Simone Martini and Andrea Masini. Experiments in linear natural deduction. Manuscript, March 1995.

[342] Simone Martini and Andrea Masini. On the fine structure of the exponential rule. In J.- Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 197–210. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[343] G. Mascari and M. Pedicini. Head linear reduction and pure proof net extraction. Theoretical Computer Science, 135:111–137, 1994.

[344] M. Masseron, C. Tollu, and J. Vauzeilles. Plan generation and linear logic. In Proceedings of the Conference on Foundations of Software Technology and Theoretical Computer Science, pages 63–75, Bangalore, India, December 1990. Springer-Verlag LNCS 472.

[345] M. Masseron, C. Tollu, and J. Vauzeilles. Planification et logique linéaire. In Proceedings of the 8ème Congrès sur la Reconnaissance des Formes et Intelligence Artificielle, pages 25–29, Lyon, France, November 1991. AFCET.

[346] M. Masseron, C. Tollu, and J. Vauzeilles. Planification et logique linéaire. Revue d’Intelligence Artificielle, pages 285–311, 1992.

[347] M. Masseron, C. Tollu, and J. Vauzeilles. Generating plans in linear logic I: Actions and proofs. Theoretical Computer Science, 113(2):349–371, 1993.

[348] M. Masseron, C. Tollu, and J. Vauzeilles. Generating plans in linear logic II: A geometry of conjunctive actions. Theoretical Computer Science, 113:371–375, 1993.

[349] F. Métayer. Homology of proof-nets. Preprint 39, Équipe de Logique Mathématique, University of Paris VII, 1992.

[350] F. Métayer. Homology of proof-nets. Archive of Mathematical Logic, 33:169–188, 1994.

[351] F. Métayer. Une étude homologique des réseaux. PhD thesis, University of Paris VII, 1994.

[352] F. Métayer. Volume of multiplicative formulas and provability. In J.-Y. Girard, Y. Lafont, and L. Regnier, editors, Advances in Linear Logic, pages 297–306. Cambridge University Press, 1995. Proceedings of the Workshop on Linear Logic, Ithaca, New York, June 1993.

[353] Daniel Mey. Investigations on a Calculus without Contractions. PhD thesis, ETH Zürich, 1992.

[354] Daniel Mey. Game-theoretical interpretation for a predicate logic without contraction. Theoretical Computer Science, 123:342–349, 1994.

[355] R. K. Meyer and R. Routley. An undecidable relevant logic. Zeitschrift für Mathematische Logik und Grundlagen der Mathematik, 19:389–397, 1973.

[356] Dale Miller. The π-calculus as a theory in linear logic: Preliminary results. In E. Lamma and P. Mello, editors, Proceedings of the Workshop on Extensions of Logic Programming, pages 242–265. Springer-Verlag LNCS 660, 1992.

[357] Dale Miller. A multiple-conclusion meta-logic. In S. Abramsky, editor, Ninth Annual Symposium on Logic in Computer Science, pages 272–281, Paris, France, July 1994. IEEE Computer Society Press.

[358] Dale Miller. A multiple-conclusion specification logic. Theoretical Computer Science, 165(1):201–232, 1996.

[359] Grigori Mints. Some information on linear logic. Manuscript, 1991.

[360] Grigori Mints. Resolution calculus for the first order linear logic. Journal of Logic, Language and Information, 2(1):59–83, 1993.

[361] E. Monteiro. Linear logic as CSP. Journal of Logic and Computation, 4(4):405–421, 1994.

[362] C. R. Murthy. A computational analysis of Girard’s translation and LC. In Seventh Annual Symposium on Logic in Computer Science, pages 90–101, Santa Cruz, California, June 1992. IEEE Computer Society Press.

[363] Sara Negri. Semantical observations on the embedding of intuitionistic logic into intuitionistic linear logic. Mathematical Structures in Computer Science, 5:41–68, 1995.

[364] S. B. Niefield and K. I. Rosenthal. Constructing locales from quantales. Mathematical Proceedings of the Cambridge Philosophical Society, 104:215–234, 1988.

[365] S. Nishizaki. Programs with continuations and linear logic. In T. Ito and A. R. Meyer, editors, Theoretical Aspects of Computer Software, Sendai, Japan, September 1991. Springer-Verlag LNCS 526.

[366] P. W. O’Hearn. Linear logic and interference control: Preliminary report. In S. Abramsky, P.-L. Curien, A. M. Pitts D. H. Pitt, , A. Poigné, and D. E. Rydeheard, editors, Proceedings of the Conference on Category Theory and Computer Science, pages 74–93, Paris, France, 1991. Springer-Verlag LNCS 530.

[367] M. Okada. Mobile linear logic as a framework for asynchronous and synchronous mobile communication calculi. Technical Report ??, Concordia University, February 1993.

[368] M. Okada. Girard’s phase semantics and a higher order cut-elimination proof. unpublished manuscript, 1994.

[369] Hiroakira Ono. Algebraic aspects of logics without structural rules. Manuscript, 1989.

[370] Hiroakira Ono. Phase structures and quantales - a semantical study of logics without structural rules. Manuscript, 1990.

[371] Hiroakira Ono. Structural rules and a logical hierarchy. In P. P. Petkov, editor, Mathematical Logic, pages 95–104. Plenum Press, 1990. Proceedings of the Summer School and Conference on Mathematical Logic, honourably dedicated to the 90th Anniversary of Arend Heyting (1898–1980), Chaika, Bulgaria, 1988.

[372] Mati Pentus. Equivalent types in Lambek calculus and linear logic. Manuscript, 1992.

[373] Mati Pentus. The conjoinability relation in Lambek calculus and linear logic. ILLC ML-93-03, University of Amsterdam, 1993.

[374] Guy Perrier. De la Construction de Preuves à la Programmation Parallèle en Logique Linéaire. PhD thesis, Université Henri Poincaré, 1995.

[375] Guy Perrier. Concurrent programming as proof net construction. Mathematical Structures in Computer Science, 00:1–50, 1997.

[376] J. F. Peters III. Timed linear logic and neural computation. Informatica y Automatica, 27:5–18, 1994.

[377] Frank Pfenning. Structural cut elimination in linear logic. Technical Report CMU-CS-94-222, Carnegie Mellon University, Department of Computer Science, December 1994.

[378] Mario Piazza. Derivability of Structural Rules in the Multiplicative Fragment of Linear Logic. PhD thesis, University of Genoa, 1995. (In Italian).

[379] Mario Piazza and M. Castellan. Quantales and structural rules. Journal of Logic and Computation, 6(5):709–724, 1996.

[380] R. Pliushkevichus. On a version of the constructive predicate calculus without structural rules. Soviet Math. Doklady, 6:416–419, 1965.

[381] Gordon D. Plotkin. Notes on event structures and Chu. Manuscript, August 1993.

[382] Gordon D. Plotkin. Notes on the Chu construction and recursion. Manuscript, August 1993.

[383] Vaughan R. Pratt. Event spaces and their linear logic. In Proceedings of the Workshop on Algebraic Methodology and Software Technology, pages 1–23, Iowa City, Iowa, 1991. Springer-Verlag.

[384] Vaughan R. Pratt. The duality of time and information. In W. Cleaveland, editor, Proceedings of the Third International Conference on Concurrent Theory, pages 237–253, Stonybrook, New York, August 1992. Springer-Verlag LNCS 630.

[385] Vaughan R. Pratt. Linear logic for generalized quantum mechanics. In Proceedings of the Workshop on Physics and Computation, pages 166–180, Dallas, Texas, October 1992. IEEE Press.

[386] Vaughan R. Pratt. The second calculus of binary relations. In A. M. Borzyszkowski and S. Sokolowski, editors, Mathematical Foundations of Computer Science, pages 142–155, Gdańsk, Poland, 1993. Springer-Verlag LNCS 781.

[387] Vaughan R. Pratt. Chu spaces: Automata with quantum aspects. In Proc. Workshop on Physics and Computation – PHYSCOMP’94, Dallas, Texas, 1994. IEEE.

[388] Vaughan R. Pratt. Chu spaces: Complementarity and uncertainty in rational mechanics. Course notes, TEMPUS summer school, 35pp, 1994.

[389] Vaughan R. Pratt. Shorter proof of universality of Chu spaces. Manuscript, August 1994.

[390] Vaughan R. Pratt. Time and information in sequential and concurrent computation. In Proceedings of the conference on Theory and Practice of Parallel Programming – TPPP’94, Sendai, Japan, November 1994.

[391] Vaughan R. Pratt. Chu spaces and their interpretation as concurrent objects. In J. van Leeuwen, editor, Computer Science Today: Recent Trends and Developments, pages 392– 405. Springer-Verlag LNCS 1000, 1995.

[392] Vaughan R. Pratt. Rational mechanics and natural mathematics. In Proceedings of the Conference on Theory and Practice of Software Development – TAPSOFT’95, pages 108– 122, Aarhus, Denmark, 1995.

[393] Vaughan R. Pratt. The stone gamut: A coordinatization of mathematics. In Tenth Annual Symposium on Logic in Computer Science, pages 444–454, San Diego, California, June 1995. IEEE Computer Society.

[394] Vaughan R. Pratt. Broadening the denotational semantics of linear logic. In Proceedings of Tokyo Conference on Linear Logic, volume 3 of Electronic Notes in Computer Science, Tokyo, Japan, 1996. Elsevier. To appear.

[395] Alberto Pravato and Luca Roversi. Lambda! considered both as a paradigmatic language and as a meta-language. Manuscript, 1995.

[396] A. Prijatelj. Investigating Bounded Contraction. PhD thesis, Institute for Logic, Language and Computation, University of Amsterdam, 1995.

[397] A. Prijatelj. Reflections on ‘difficult’ embeddings. Journal of Philosophical Logic, 24:71–84, 1995.

[398] A. Prijatelj and Harold Schellinx. A note on embeddings of intuitionistic logic into classical linear logic. Manuscript, July 1991.

[399] M. Quatrini. Sémantique cohérente des exponentielles: de la logique linéaire à la logique classique. PhD thesis, University of Aix-Marseille II, 1995.

[400] Udday S. Reddy. Acceptors as values: Functional programming in classical linear logic. Manuscript, December 1991.

[401] Udday S. Reddy. Linear logic model of state. Manuscript, October 1993.

[402] Udday S. Reddy. A typed foundation for directional logic programming. In E. Lamma and P. Mello, editors, Third International Workshop on Extensions of logic programming, pages 150–167, Bologna, Italy, 1993. Springer-Verlag LNAI 660.

[403] Laurent Regnier. Lambda-calculus and its dynamic algebra II: Computations. Manuscript, July 1989.

[404] Laurent Regnier. λ-Calcul et Réseaux. PhD thesis, University of Paris VII, 1992.

[405] Christian Retoré. A first move from non-commutative linear logic to true concurrency: the calculus of ordered sequents. Manuscript, 1990.

[406] Christian Retoré. Graph theory from linear logic: Aggregates. Preprint 47, Équipe de Logique Mathématique, University of Paris VII, October 1993.

[407] Christian Retoré. Réseaux and Séquents Ordonnés. PhD thesis, University of Paris VII, 1993.

[408] Christian Retoré. On the relation between coherence semantics and multiplicative proof nets. Manuscript, 1994.

[409] Simona Ronchi della Rocca and Luca Roversi. Lambda calculus and intuitionistic linear logic. In Logic Colloquium ’94, Clermont Ferrand, France, 1994.

[410] Dirk Roorda. Investigations into classical linear logic. ITLI-Prepublication Series for Mathematical Logic and Foundations ML-89-06, University of Amsterdam, 1989.

[411] Dirk Roorda. Quantum graphs and proof nets. Manuscript, 1990.

[412] Dirk Roorda. Resource Logics: Proof-theoretical Investigations. PhD thesis, University of Amsterdam, September 1991.

[413] Dirk Roorda. Proof nets for Lambek calculus. Journal of Logic and Computation, 2:211– 231, 1992.

[414] Dirk Roorda. Interpolation in fragments of classical linear logic. Journal of Symbolic Logic, 59(2):419–444, 1994.

[415] K. I. Rosenthal. A note on Girard’s quantales. Cahiers de Géométry différentielle Catégorique, 31:1–11, 1990.

[416] K. I. Rosenthal. Quantales and their Applications. Longman Scientific & Technical, 1990. Pitman Research Notes in Mathematics Series 234.

[417] K. I. Rosenthal. Girard quantaloids. Mathematical Structures in Computer Science, 2(1):93–108, 1992.

[418] Luca Roversi. A compiler from Curry-typed λ-terms to the linear λ-terms. In P. Mentrasti and M. Venturini Zilli, editors, Proceedings of the Fourth Italian Conference Theoretical Computer Science, pages 330–344, L’aquila, Italy, 1992. World Scientific Publishing Company.

[419] Paul Ruet. Non-commutative linear logic with mobilities. In Proceedings of Logic Colloquium ’96, pages 274–275, June 1996.

[420] Paul Ruet. Non-Commutative Logic and Concurrent Constraint Programming. PhD thesis, Université Denis Diderot, Paris 7, 1997.

[421] Paul Ruet. Phase semantics for mixed non-commutative classical linear logic. Technical report, LIENS, 1997.

[422] Colin Runciman and David Wakeling. Linearity and laziness. In J. Hughes, editor, Proceedings of the Fifth ACM Conference on Functional Programming Languages and Computer Architecture, pages 215–240, Cambridge, Massachusetts, 1991. Springer-Verlag LNCS 666.

[423] Giovanni Sambin. Intuitionistic formal spaces and their neighbourhood. In C. Bonotto, R. Ferro, S. Valentini, and A. Zanardo, editors, Logic Colloquium ’88, pages 261–286. North-Holland, 1989.

[424] Vijay Saraswat. A brief introduction to linear concurrent constraint programming. Manuscript, 1993.

[425] Andre Scedrov. A brief guide to linear logic. In G. Rozenberg and A. Salomaa, editors, Current Trends in Theoretical Computer Science, pages 377–394. World Scientific Publishing Company, 1993. Also in Bulletin of the European Association for Theoretical Computer Science, volume 41, pages 154–165.

[426] Andre Scedrov. Linear logic and computation: A survey. In L. F. Bauer, W. Brauer, and H. Schwichtenberg, editors, Proceedings of the International Summer School of Marktoberdorf, NATO Advanced Science Institutes, series F94, pages ??–?? Springer-Verlag, 1994.

[427] Harold Schellinx. Introducing: “Linear … Logic …!?!”. Manuscript, September 1991.

[428] Harold Schellinx. Some syntactical observations on linear logic. Journal of Logic and Computation, 1(4):537–559, 1991.

[429] Harold Schellinx. A linear approach to modal proof theory. In Proceedings of a Workshop on the Proof Theory of Modal Logic, pages ??–??, Hamburg, Germany, March 1993.

[430] Harold Schellinx. A hitchhiker’s guide to linear decorating. Manuscript, 1994.

[431] Harold Schellinx. The Noble Art of Linear Decorating. PhD thesis, Institute for Logic, Language and Computation, University of Amsterdam, 1994.

[432] P. Schröder-Heister. Structural frameworks, substructural logics, and the role of elimination inferences. In G. Huet and G. Plotkin, editors, Logical Frameworks, pages 385–403. Cambridge University Press, 1991.

[433] R. A. G. Seely. Linear logic, ⋆-autonomous categories and cofree coalgebras. In J. W. Gray and A. Scedrov, editors, Categories in Computer Science and Logic, pages 371–382. American Mathematical Society, 1989. Proceedings of the AMS-IMS-SIAM Joint Summer Research Conference, June 14–20, 1987, Boulder, Colorado; Contemporary Mathematics Volume 92.

[434] R. A. G. Seely. Polymorphic linear logic and topos models. C.R. Math. Rep. Acad. Sci. Canada - Vol XII, No. 1, February 1990.

[435] M. Shiharata. Linear Set Theory. PhD thesis, Stanford University, 1994.

[436] U. Solitro. A typed calculus based on a fragment of linear logic. Theoretical Computer Science, 68:333–342, 1989.

[437] U. Solitro. La Logica Lineare come Sistema di Tipi per un Calcolo non Funzionale. PhD thesis, Dipartimento di Scienze dell’Informazione, University of Milano, 1990.

[438] U. Solitro and S. Valentini. Local computations in linear logic. Mathematical Logic, 2:??– ??, 1993.

[439] T. Tammet. Proof strategies in linear logic. Journal of Automated Reasoning, 12:273–304, 1994. Also available as Programming Methodology Group Report 70, Chalmers University, 1993.

[440] P. Taylor. Quantitative domains, groupoids and linear logic. In P. Dybjer, A. M. Pitts, D. H. Pitt, A. Poigné, and D. E. Rydeheard, editors, Proceedings of the Conference on Category Theory and Computer Science, pages 155–181. Springer-Verlag LNCS 389, September 1989. Manchester, UK.

[441] C. Tollu. Planification en Logique Linéaire et Langage de Requête Relationelles avec Compteur. PhD thesis, Paris, 1993.

[442] L. Tortora de Falco. Strong normalization property for heterostyle LKtq and FD. Preprint ??, Équipe de Logique Mathématique, University of Paris VII, 1995.

[443] T. H. Trimble. Linear Logic, Bimodules, and Full Coherence for Autonomous Categories. PhD thesis, Rutgers University, 1994.

[444] Anne S. Troelstra. Lectures on Linear Logic. CSLI Lecture Notes 29, Center for the Study of Language and Information, Stanford, California, 1992.

[445] Anne S. Troelstra. Natural deduction for intuitionistic linear logic. Prepublication Series for Mathematical Logic and Foundations ML-93-09, Institute for Language, Logic and Computation, University of Amsterdam, 1993.

[446] Alasdair Urquhart. Semantics for relevant logic. Journal of Symbolic Logic, 37:159–169, 1972.

[447] Alasdair Urquhart. The undecidability of entailment and relevant implication. Journal of Symbolic Logic, 49:1059–1073, 1984.

[448] Alasdair Urquhart. The complexity of decision procedures in relevance logic. Technical Report 217/89, Department of Computer Science, University of Toronto, 1989.

[449] Alasdair Urquhart. The complexity of linear logic. Manuscript, March 1990.

[450] Aldo Ursini. Semantical investigations of linear logic. Manuscript, December 1995.

[451] J. Vauzeilles. Cut elimination for the unified logic. Annals of Pure and Applied Logic, 62:1–6, 1993.

[452] Paolo Volpe. Concurrent logic programming as uniform linear proofs. In G. Levi and M. Rodríguez-Artalejo, editors, Proceedings of the Fourth International Conference on Algebraic and Logic Programming, pages 133–149, Madrid, Spain, September 1994. Springer-Verlag LNCS 850.

[453] Philip Wadler. Linear types can change the world. In M. Broy and C. B. Jones, editors, IFIP TC 2 Working Conference on Programming Concepts and Methods, pages 561–581, Sea of Gallilee, Israel, April 1990. North-Holland.

[454] Philip Wadler. Is there a use for linear logic? In Proceedings of the Symposium on Partial Evaluations and Semantics-Based Program Manipulation, pages 255–273, New Haven, Connecticut, June 1991. Also in SIGPLAN Notices; vol.26, no.9; September 1991.

[455] Philip Wadler. There’s no substitute for linear logic. Manuscript, December 1991.

[456] Philip Wadler. A syntax for linear logic. In S. Brookes, M. Main, A. Melton, M. Mislove, and D. Schmidt, editors, Proceedings of the Ninth Conference on Mathematical Foundations of Programming Semantics, pages 513–529, New Orleans, Louisiana, April 1993. Springer-Verlag LNCS 647.

[457] Philip Wadler. A taste of linear logic. In A. M. Borzyszkowski and S. Sokolowski, editors, Mathematical Foundations of Computer Science, pages 185–210, Gdańsk, Poland, 1993. Springer-Verlag LNCS 781.

[458] David Wakeling. Linearity and Laziness. PhD thesis, Department of Computer Science, University of York, November 1990. Available as Technical Report YCST 90/07.

[459] Michael Winikoff. Hitch hiker’s guide to Lygon 0.7. Technical Report TR 96/36, Melbourne University, Department of Computer Science, 1996.

[460] Michael Winikoff and James Harland. Determining logic programming languages. Technical Report TR 95/26, Melbourne University, Department of Computer Science, 1995.

[461] David N. Yetter. Quantales and (noncommutative) linear logic. Journal of Symbolic Logic, 55(1):41–64, March 1990.

[462] J. Zlatuška. Committed-choice concurrent logic programming in linear logic. In G. Gottlob, A. Leitsch, and D. Mundici, editors, Computational Logic and Proof Theory, Third Kurt Gödel Colloquium, pages 337–348, Brno, Czech Republic, August 1993. Springer-Verlag LNCS 713.

[463] J. Zlatuška. Linear logic semantics for concurrent prolog. In R. Dyckhoff, editor, Fourth International Workshop on Extensions of Logic Programming, pages 348–360, St. Andrews, Scotland, March 1993. Springer-Verlag LNAI 798.